OpenStudy (anonymous):

At -3ºC vapor pressure of ice is 3.566mm and for supercooled water is 3.669mm. For a transition of 1 mol of supercooled water into 1 mol of ice at -3ºC and 3.669 mm, compute delta G. (Ice density = 0.917 g/cm^3) (Solution: -63.9 J)

OpenStudy (anonymous):

@.Sam. ??

OpenStudy (anonymous):

do u know the formula \[dg=-s \ dT+ v \ dP\] ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

well T is constant for this transition so dT=0 and formula becomes to \[dg=v \ dP\]

OpenStudy (anonymous):

Yes, I've done that but I don't get -63.9 J (the solution in the book), maybe I'm wrong with my calculus but I dont' know where

OpenStudy (anonymous):

sorry i lost my connection let me check it

OpenStudy (anonymous):

Ok, np, thank you!

OpenStudy (anonymous):

what was your answer?

OpenStudy (anonymous):

A non-sense I thinkk, 2.02 J

OpenStudy (anonymous):

\[\Delta P=-0.113 \ \ mmHg\] am i right?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

and we assume that v is constant ok?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[v=\frac{1}{0.917}\frac{cm^3}{gr}18\frac{gr}{mol}=19.63\frac{cm^3}{mol}\]

OpenStudy (anonymous):

\[\Delta P=-0.113 \ \ mmHg=-0.113 \frac{101325}{760}=15.065 \ \ Pa\]

OpenStudy (anonymous):

-15.065*

OpenStudy (anonymous):

my answer is a very small number o.O

OpenStudy (anonymous):

-0.002 J

OpenStudy (anonymous):

what shall we do?

OpenStudy (anonymous):

And what about using Clausius-Clapeyron equation? We can find out latent heat's

OpenStudy (anonymous):

but this transition starts with a supercooled water for which we dont have any information about its state

OpenStudy (anonymous):

\[\Delta G=-SdT+VdP\] as T=constant \[\Delta G =VdP\] \[\Delta G=\int\limits_{3.669}^{3.566} VdP = RTln(3.566/3.669)=8.31(-3+273.15)\ln(3.566/3.669)=-63.9 (J)\]

OpenStudy (anonymous):

If you considered water vapor as an ideal gas you can make this substitution

OpenStudy (anonymous):

Anyway thanks a lot @mukushla! ^^

OpenStudy (anonymous):

thats ok thank u for letting me know the correct answer